OpenStudy (anonymous):

simplify(7-√3)(-6+√3)

OpenStudy (anonymous):

= -42 - 3 +13 sqrt(3) = -45 +13sqrt(3)

OpenStudy (anonymous):

\[(7-\sqrt{3})(-6+\sqrt{3})=-42-3+7\sqrt{3}+6\sqrt{3}\]\] \[=-45+13\sqrt{3}\]